Jacquie Inc. reports the following annual cost data for its single product. Normal production and sales level 60,000 units Sales price 56.00 per unit Direct materials 9.00 per unit Direct labor 6.50 per unit Variable overhead 11.00 per unit $720,000 in total Fixed overhead Complete the below table using absorption costing. (Round cost per unit answers to 2 decimal place.) Production volume Cost of goods sold 60,000 units 80,000 units Cost of goods sold per unit Number of units sold Total cost of goods sold
Answer #1
Production Volume Working Notes
Cost of goods sold 60000 units 80000 units Total Fixed overhead will remain unchanged regardless of level of activity
Direct Material 9 9 i.e. 720000/60000 units= 12 per unit and
Direct Labor 6.5 6.5 720000/80000 units= 9 per unit
Variable Overhead 11 11
Fixed overhead 12 9
Cost of goods sold per units      A 38.5 35.5
Number of units sold              B 60000 80000
Total cost of goods sold        A*B 2310000 2840000
Jacquie Inc
Income statement through gross margin
Sales Volume
60000 units 80000 units
Sales Revenue(per unit i.e. $56 * total units) 3360000 4480000
Cost of goods sold from above table 2310000 2840000
Gross Margin 1050000 1640000
If jacquie increases its production to 80000 units, while sales remain at the
current 60000 unit level, by how much would the company's gross margin
increase or decrease under absorption costing? Assume company has idle
capacity to double current production.
Number of units sold 60000 Working Notes
Change in fixed overhead cost per unit 3 change in fixed overhead cost per unit =12-9   = 3
change in cost of goods sold 180000
For your understanding & reconciliation purpose only
Cost sheet under absorption costing units per unit
Production Volume 80000
Direct Material 9 720000
Direct Labor 6.5 520000
Variable Overhead 11 880000
Thank You Fixed overhead 720000
Cost of manufacturing/production 2840000
Less : Closing stock of finished goods 20000 35.5 710000
Cost of goods sold with Inventory 2130000
Cost of goods sold without Inventory 2310000
Change in cost of goods sold 180000